# TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design

### v2

@@ F.1 Minigrid Environment
Figure 13: MiniGrid zero-shot environments used for evaluation. In SmallCorridor and LargeCorridor, the goal can appear in any of the corridor paths. SimpleCrossing and FourRooms are adapted from MiniGrid [16], and PerfectMaze from REPAIRED [6].
@@
The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid [16]. Each tile in the maze can be either an empty tile, a wall, a goal, or the agent itself. The empty tile is a navigable space through which the agent can move, whereas the wall is an impassable obstacle that blocks the agent’s movement. At the beginning of each episode, the initial position of the agent, the goal location, and the wall layout are randomly initialized. Up to 60 walls can be placed throughout the maze, increasing navigational complexity. At each timestep, the agent receives a $5\times 5$ local observation, along with its current direction. The action space is defined as a discrete set of seven possible actions, although only three—turn left, turn right, and move forward—are relevant to maze navigation. These actions are selected without masking the unused action outputs. A reward is provided when the agent successfully reaches the goal, calculated as $1-T/T_{\text{max}}$, where $T$ is the number of steps taken and $T_{\text{max}}=250$ denotes the maximum allowed steps per episode. If the agent fails to reach the goal within the time limit, it does not receive a reward.
@@
To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work [4, 6]. Figure 13 displays all evaluation tasks.
F.2 BipedalWalker Environment
@@
The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym [23]. At each timestep, the agent receives a 24-dimensional observation, which includes information such as the hull angle and angular velocity, joint angles and speeds for the hips and knees, ground contact indicators for each foot, the robot’s horizontal and vertical velocities, and lidar-based distance measurements to the ground ahead. The action space is represented by a 4-dimensional vector that controls the torques applied to the robot’s two hip joints and two knee joints. Rewards are structured to encourage efficient and stable locomotion across the terrain. The agent receives positive rewards for processing while maintaining a straight hull posture, with small penalties applied for energy expenditure to discourage unnecessary motor use. The maximum achievable score is $300$. If the agent falls or moves backward, the episode ends immediately, and a penalty of $-100$ is applied. The maximum episode length is 2000 timesteps. To introduce additional difficulty and test the robustness of the agent’s locomotion policy, the environment includes a variety of challenging terrain features. These include stairs, which consist of sequences of elevated steps that the agent must ascend or descend; pit gaps, which are horizontal gaps over which the agent must jump or step over; stumps, which are vertical obstacles with varying heights and widths that obstruct movement; and surface roughness, which introduces uneven ground textures that can disrupt balance and foot placement.
@@
We evaluate agents in the BipedalWalker-v3 environment, which features relatively smooth terrain, as well as in the more demanding BipedalWalkerHardcore-v3, which introduces a variety of challenging obstacles. The BipedalWalker-v3 focuses on basic locomotion and balance without significant terrain disturbances. Terrain in this environment does not have obstacles such as stairs, pit gaps, and stumps, and only gentle slopes are generated by adding noise in the range of approximately -1 to 1. In contrast, BipedalWalkerHardcore-v3 includes complex terrain elements such as stairs, pit gaps, stumps, and roughness, which require precise control and adaptive strategies. These obstacles are procedurally generated with randomized parameters. The stair height is randomly set to either $+1$ (ascending) or $-1$ (descending), with the number of steps and step width randomly sampled between 3–5 steps and 4–5 units, respectively. The pit width is determined by sampling a single integer between 3–5 units. Both the height and width of each stump are sampled as a single random integer between 1–3 units, ensuring the stumps are always square-shaped. Terrain roughness noise range is the same as BipedalWalker-v3.
@@
Figure 14: Evaluation Terrains in BipedalWalker. Six held-out environments used to assess zero-shot generalization: (a) BipedalWalker-v3 with gentle noise-induced slopes; (b) BipedalWalkerHardcore-v3 combining multiple obstacle types; (c) Stairs, a fixed 5-step sequence requiring precise stepping; (d) PitGap, a 5-unit horizontal gap; (e) Stump, 2-unit tall obstacles with variable width; and (f) Roughness, continuous uneven ground generated by noise.
F.3 Network Structures
For partially observable navigation, the LSTM-based transition prediction model $f_{\phi}$ consists of three components: 1. an image encoder that compresses the input observation $s_{t}$ into a 128-dimensional latent embedding, 2. an LSTM module (hidden size 128, 1 layer) that processes the concatenation of this embedding and the action $a_{t}$, 3. and an image decoder that reconstructs the predicted next observation $\hat{s}_{t+1}$ from the LSTM output. We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video-prediction tasks under partial observability [25, 26].
@@ F.3 Network Structures
We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction [27], while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains [28].
The PPO network serves as the student agent in our system, directly learning to optimize both the policy and value functions through interaction with the environment. The architecture follows the standard actor-critic structure where the actor head outputs an action distribution, and the critic head predicts the scalar value $V(s)$.
@@
Appendix G Overall Workflow
Algorithm 1 summarizes our overall UED procedure under the TRACED framework. At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
If the buffer is full, we evict the lowest-priority task before appending. Otherwise, with probability $p_{\mathrm{replay}}$ the teacher enters the replay+mutation phase, drawing a batch of $B$ tasks from the buffer proportional to their TaskPriority scores, updating the policy on each, and recomputing regret. We then select the $N_{\mathrm{mutate}}$ tasks with the smallest regret, mutate each via the editor, evaluate the new variant, and replace its parent in the buffer. Finally, for every modified task we recompute TaskDifficulty, CoLearnability, and TaskPriority (Eq. (7)) before proceeding to the next timestep.
Algorithm 1 UED Workflow with Transition-Prediction Loss and Co-Learnability
3:while policy has not converged do
4:     Sample decision $d_{t}\sim\mathrm{Bernoulli}(p_{\mathrm{replay}})$
5:     if $d_{t}=0$ $\triangleright$ Exploration phase
6:         Sample new task $\theta\sim\mathcal{G}$
8:         Compute regret estimate
9:         Append $\theta$ to buffer $\Lambda$ (if $|\Lambda|>K$, remove lowest-priority)
10:     $\triangleright$ Replay + Mutation phase
11:         Sample a batch $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ w.p. $\propto$ $\mathrm{TaskPriority}(\cdot,t)$
12:         for $k=1\ldots B$ do
13:              Collect trajectory $\tau_{k}$ by executing $\pi_{\phi}$ on $\theta_{k}$
14:              Update policy $\phi$ using rewards from $\tau_{k}$
15:              Compute $\widehat{\mathrm{Regret}}(\tau_{k})$
16:         end for
18:         for each selected $\theta_{k^{\prime}}$ do
19:              $\tilde{\theta}\leftarrow$ editor($\theta_{k^{\prime}}$)
20:              Collect trajectory $\tilde{\tau}$ on $\tilde{\theta}$ (no gradient)
21:              Compute $\widehat{\mathrm{Regret}}(\tilde{\tau})$
22:              Replace $\theta_{k^{\prime}}$ in $\Lambda$ with $\tilde{\theta}$
23:         end for
24:     end if
25:     Buffer update:
26:     for each task $i\in\Lambda$ updated at $t$ do
27:         Recompute $\mathrm{TaskDifficulty}(i,t)$
28:         Recompute $\mathrm{CoLearnability}(i,t-1)$
29:         Update $\mathrm{TaskPriority}(i,t)$ via Eq. (7)
30:     end for
31:     $t\leftarrow t+1$
32:end while
Appendix H Hyperparameters
Table 13 contains the hyperparameters we use. Hyperparameter values were adopted directly from the DCD repository [29]. We perform hyperparameter tuning in MiniGrid and Bipedal environments, focusing on two key components, $\alpha$ and $\beta$. We explored values in the ranges $\alpha\in[0.5,1.0,1.5]$, $\beta\in[0.6,0.8,1.0]$.

### v3

@@ H.1 Minigrid Environment
Figure 20: MiniGrid zero-shot environments used for evaluation. In SmallCorridor and LargeCorridor, the goal can appear in any of the corridor paths. SimpleCrossing and FourRooms are adapted from MiniGrid (Chevalier-Boisvert et al., 2023), and PerfectMaze from REPAIRED (Jiang et al., 2021a).
@@
The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid (Chevalier-Boisvert et al., 2023). Each tile in the maze can be either an empty tile, a wall, a goal, or the agent itself. The empty tile is a navigable space through which the agent can move, whereas the wall is an impassable obstacle that blocks the agent’s movement. At the beginning of each episode, the initial position of the agent, the goal location, and the wall layout are randomly initialized. Up to 60 walls can be placed throughout the maze, increasing navigational complexity. At each timestep, the agent receives a $5\times 5$ local observation, along with its current direction. The action space is defined as a discrete set of seven possible actions, although only three, turn left, turn right, and move forward, are relevant to maze navigation. These actions are selected without masking the unused action outputs. A reward is provided when the agent successfully reaches the goal, calculated as $1-T/T_{\text{max}}$, where $T$ is the number of steps taken and $T_{\text{max}}=250$ denotes the maximum allowed steps per episode. If the agent fails to reach the goal within the time limit, it does not receive a reward.
@@
To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work (Dennis et al., 2020; Jiang et al., 2021a). Figure 20 displays all evaluation tasks.
H.2 BipedalWalker Environment
@@
The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym (Brockman et al., 2016). At each timestep, the agent receives a 24-dimensional observation, which includes information such as the hull angle and angular velocity, joint angles and speeds for the hips and knees, ground contact indicators for each foot, the robot’s horizontal and vertical velocities, and lidar-based distance measurements to the ground ahead. The action space is represented by a 4-dimensional vector that controls the torques applied to the robot’s two hip joints and two knee joints. Rewards are structured to encourage efficient and stable locomotion across the terrain. The agent receives positive rewards for processing while maintaining a straight hull posture, with small penalties applied for energy expenditure to discourage unnecessary motor use. The maximum achievable score is $300$. If the agent falls or moves backward, the episode ends immediately, and a penalty of -100 is applied. The maximum episode length is 2000 timesteps. To introduce additional difficulty and test the robustness of the agent’s locomotion policy, the environment includes a variety of challenging terrain features. These include stairs, which consist of sequences of elevated steps that the agent must ascend or descend; pit gaps, which are horizontal gaps over which the agent must jump or step over; stumps, which are vertical obstacles with varying heights and widths that obstruct movement; and surface roughness, which introduces uneven ground textures that can disrupt balance and foot placement.
@@
We evaluate agents in the BipedalWalker-v3 environment, which features relatively smooth terrain, as well as in the more demanding BipedalWalkerHardcore-v3, which introduces a variety of challenging obstacles. The BipedalWalker-v3 focuses on basic locomotion and balance without significant terrain disturbances. Terrain in this environment does not have obstacles such as stairs, pit gaps, and stumps, and only gentle slopes are generated by adding noise in the range of approximately -1 to 1. In contrast, BipedalWalkerHardcore-v3 includes complex terrain elements such as stairs, pit gaps, stumps, and roughness, which require precise control and adaptive strategies. These obstacles are procedurally generated with randomized parameters. The stair height is randomly set to either +1+1 (ascending) or 1-1 (descending), with the number of steps and step width randomly sampled between 3-5 steps and 4-5 units, respectively. The pit width is determined by sampling a single integer between 3-5 units. Both the height and width of each stump are sampled as a single random integer between 1-3 units, ensuring the stumps are always square-shaped. Terrain roughness noise range is the same as BipedalWalker-v3.
@@
Figure 21: Evaluation Terrains in BipedalWalker. 6 held-out environments used to assess zero-shot generalization: (a) BipedalWalker-v3 with gentle noise-induced slopes; (b) BipedalWalkerHardcore-v3 combining multiple obstacle types; (c) Stairs, a fixed 5-step sequence requiring precise stepping; (d) PitGap, a 5-unit horizontal gap; (e) Stump, 2-unit tall obstacles with variable width; and (f) Roughness, continuous uneven ground generated by noise.
H.3 Network Structures
For partially observable navigation, the LSTM-based transition prediction model $f_{\phi}$ consists of three components: 1. an image encoder that compresses the input observation $s_{t}$ into a 128‑dimensional latent embedding, 2. an LSTM module (hidden size 128, 1 layer) that processes the concatenation of this embedding and the action $a_{t}$, 3. and an image decoder that reconstructs the predicted next observation $\hat{s}_{t+1}$ from the LSTM output. We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video‑prediction tasks under partial observability (Ha & Schmidhuber, 2018; Oh et al., 2015).
@@ H.3 Network Structures
We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction (Zhao et al., 2016), while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains (Mnih et al., 2015).
We adopt ACCEL’s network architecture (Parker-Holder et al., 2022), differing only in the LSTM design choice. The PPO network serves as the student agent in our system, directly learning to optimize both the policy and value functions through interaction with the environment. The architecture follows the standard actor-critic structure where the actor head outputs an action distribution, and the critic head predicts the scalar value $V(s)$.
@@
Appendix I Difference with Former Studies
I.1 Difference with ACCEL
I.2 Difference with CENIE
CENIE introduces a complementary curriculum strategy by explicitly measuring environmental novelty, rather than relying solely on regret signals as in PLR or ACCEL. It fits a Gaussian mixture model (GMM) to the agent’s past trajectories and scores new tasks by how poorly the GMM explains them, thereby steering the agent toward unfamiliar regions of the state-action space. This model-agnostic approach can be applied on top of any UED framework, including PLR and ACCEL.
Appendix J Related Works
Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments (Wang et al., 2019; 2020; Dennis et al., 2020; Chung et al., 2024). PAIRED (Dennis et al., 2020) trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments. However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula. To mitigate these limitations, Mediratta et al. (2023) introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr (Azad et al., 2023a) employs variational autoencoder-based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
Another line of work is based on Prioritized Level Replay (PLR) (Jiang et al., 2021b), which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks. This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent. PLR  (Jiang et al., 2021a) make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer. This leads to better performance on unseen environments counterintuitively. ACCEL (Parker-Holder et al., 2022) extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer. This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks. However, these methods (Jiang et al., 2021b; a; Parker-Holder et al., 2022) approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
Moving beyond regret-based approaches, CENIE (Teoh et al., 2024) integrates novelty with regret to expose agents to more diverse learning situations. Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer. This curriculum-awareness leads to better exploration and generalization. DIPLR (Li et al., 2023) introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state-action trajectories. MBeDED (Li et al., 2024) introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task. It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task. So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains. For Multi-Agent Reinforcement Learning, MAESTRO (Samvelyan et al., 2023) extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations. In contrast, RACCOON (Erlebach & Cook, 2024) applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration. For meta-RL, GROOVE (Jackson et al., 2023) suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training. DIVA (Costales & Nikolaidis, 2024) is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage. In sim-to-real settings, GCL (Wang et al., 2024) aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task (Narvekar et al., 2020). Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress (Klink et al., 2020). Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process (Eimer et al., 2021; Klink et al., 2022). WAKER (Rigter et al., 2024) leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios. However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
Appendix K More Details On Baseline Algorithms
Robust PLR (PLR(Jiang et al., 2021a) was introduced as an extension of the original PLR (Jiang et al., 2021b). While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay. When sampling new tasks, the agent’s policy parameters are stop-gradiented, meaning the policy is not updated on trajectories collected from these newly sampled tasks. The collected trajectories are used to compute PVL scores to decide whether to add the task to the replay buffer. Policy updates are exclusively performed on tasks sampled from the replay buffer. By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
Adversarially Compounding Complexity by Editing Levels (ACCEL) (Parker-Holder et al., 2022) actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space. This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret. ACCEL maintains a buffer $\Lambda$ and employs a cycle of sampling, editing, and curation. The key insight of ACCEL is that regret serves as a domain-agnostic fitness function for evolution, enabling it to produce batches of levels at the frontier of agent capabilities. The editing mechanism involves making small mutations to previously high-regret levels, which can operate directly on environment elements such as blocks in MiniGrid. This evolutionary process creates an expanding frontier that matches the agent’s capabilities, starting with simple levels and progressively increasing in complexity. Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
Coverage-based Evaluation of Novelty In Environment (CENIE) (Teoh et al., 2024) introduces environment novelty as a complementary objective to regret in UED. CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum. The framework operates on the intuition that a novel environment should induce unfamiliar experiences, pushing the student agent into unexplored regions of the state-action space. At the core of CENIE’s implementation is the use of Gaussian Mixture Models (GMMs) to model the distribution of state-action pairs from the agent’s past experiences. Given a state-action buffer $\Gamma$ containing pairs collected from previous environments, CENIE fits a GMM with parameters $\lambda_{\Gamma}$ to represent this distribution. The novelty of a candidate environment is then quantified by measuring the similarity between its newly observed state-action pairs and the learned distribution of past state-action experiences. CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) (Chung et al., 2024) frames environment generation as a diffusion process steered by the agent’s regret. Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase (Jiang et al., 2022), and assessed ADD with the ADD codebase (Chung et al., 2024). CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
Appendix L Limitations
Co-Learnability. In this paper, we proposed Co-Learnability, simple yet effective method. Follow-up studies could address these gaps by developing multi-step Co-Learnability estimators, incorporating importance-weighting or counterfactual techniques to reduce sampling bias, and combining observational co-learn signals with auxiliary models (e.g., task-conditioned value predictors) for stronger causal inference.
ATPL. In environments with highly stochastic transitions, the ATPL signal (transition-prediction error) can become noisy, potentially reducing the fidelity of our regret approximation.
Sampling levels. We sampled a fixed number of the lowest-regret levels as easy tasks to mirror ACCEL’s approach and enable fair comparisons with baseline methods. However, this simple strategy leaves room to explore more advanced task-selection mechanisms.
Heuristic weighting. We introduced weighted approach to regret approximation and task prioritization that relies on two key parameters $\alpha$, PVL and ATPL in the regret approximation, and $\beta$, which weights the relative importance of Co-Learnability against task difficulty. While these parameters were manually tuned in the current implementation and demonstrated effective results across environments, future work could explore more principled methods for automatically determining these weights.
Extension to other RL algorithms. Our experiments have been conducted with PPO-based student agents. While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum. Off-policy methods such as SAC (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018) feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework. Moreover, Model-based approaches like DreamerV3 (Hafner et al., 2025) build world models for planning, offering complementary insights into curriculum design. Evaluating these methods under our curriculum framework remains an important direction for future work.
Extend to other UED methods. We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost. While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR (Jiang et al., 2021a) and CENIE (Teoh et al., 2024)) and conduct comprehensive experiments to assess its performance in these settings.
Evaluation in other domains. Our experimental validation has been conducted in MiniGrid (Chevalier-Boisvert et al., 2023; Dennis et al., 2020) and BipedalWalker (Brockman et al., 2016; Parker-Holder et al., 2022). Evaluating TRACED in other domains such as Car Racing (Brockman et al., 2016; Jiang et al., 2021a), physical robotics applications (Mahmood et al., 2018), and reasoning task ARC (Chollet, 2019) would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
Appendix M Experimental Setup and Reproducibility
Appendix N LLM Usage
We used a large language model (LLM) solely for language editing. Concretely, the LLM assisted with grammar and style polishing, LaTeX phrasing (e.g., equation and caption wording), and improving clarity and concision of author-written text. The LLM was not used to generate ideas, design algorithms, select hyperparameters, run experiments, analyze data, create figures/tables, write code, or produce mathematical results.
Appendix O Hyperparameters
Table 10 summarizes the hyperparameters used. Unless otherwise noted, values are adopted directly from the DCD repository (Jiang et al., 2022). The only deviation from the DCD defaults is the number of PPO workers (see Appendix A.4). We tune two key weights, $\alpha$ and $\beta$, on MiniGrid and BipedalWalker, exploring $\alpha\in\{0.5,1.0,1.5\}$ and $\beta\in\{0.6,0.8,1.0\}$.